# A Composite-Loss Graph Neural Network for the Multivariate Post-Processing of Ensemble Weather Forecasts

### v1

@@ Abstract
The aim of this study is the multivariate post-processing of ensemble forecasts using a graph neural network (dualGNN) trained with a composite loss function that combines the energy score (ES) and the variogram score (VS). The method is evaluated on two datasets: WRF-based solar irradiance forecasts over northern Chile and ECMWF visibility forecasts for Central Europe. The dualGNN consistently outperforms all empirical copula-based post-processed forecasts and shows significant improvements compared to graph neural networks trained solely on either the continuous ranked probability score (CRPS) or the ES, according to the evaluated multivariate verification metrics. Furthermore, for the WRF forecasts, the rank-order structure of the dualGNN forecasts captures valuable dependency information, enabling a more effective restoration of spatial relationships than either the raw numerical weather prediction ensemble or historical observational rank structures. By contrast, for the visibility forecasts, the GNNs trained on CRPS, ES, or the ES-VS combination outperform the calibrated reference.
@@ 3.2 Univariate post-processing of visibility
$P\!\big{(}Y\leq y_{i}\mid\mathbf{x}\big{)}=\frac{\exp\!\big{(}L_{i}(\mathbf{x})\big{)}}{1+\exp\!\big{(}L_{i}(\mathbf{x})\big{)}},\qquad L_{k}(\mathbf{x}):=\alpha_{i}+\mathbf{x}^{\top}\boldsymbol{\beta},\qquad i=1,2,\ldots,84,$
where $\alpha_{k}\in\mathbb{R}$ and $\boldsymbol{\beta}\in\mathbb{R}^{M}$ are model parameters satisfying $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{84}$. Consequently, fitting a POLR model to the visibility data requires estimating $84+M$ parameters.
@@ 3.4 Verification metrics
For a predictive cumulative distribution function $F$ and an observation $y\in\mathbb{R}$, the CRPS is defined as
@@
Moreover, for an ensemble forecast $\{f_{1},\dots,f_{K}\}$ the above reduces to the sample CRPS, that is
@@
If $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are independent $D$-dimensional random vectors with distribution $F$, the generalization of Equation (3.1) leads to the sample energy score (ES; Gneiting and Raftery,, 2007). For an ensemble forecast $\mathbf{f}_{k}=\big{(}f_{k}^{(1)},f_{k}^{(2)},\dots,f_{k}^{(D)}\big{)}^{\top}$, $k=1,\dots,K$, and the corresponding observation vector $\mathbf{y}=\big{(}y^{(1)},y^{(2)},\dots,y^{(D)}\big{)}^{\top}$, the sample energy score is given by
@@
Graphical tools for assessing the reliability of multivariate forecasts, such as multivariate rank histograms, are also well established. These histograms are based on pre-ranks, which condense a multivariate quantity into a single value. Since interpreting multivariate histograms can be challenging, it is advisable to employ multiple pre-ranking methods concurrently to obtain a comprehensive view of forecast miscalibration (Allen et al.,, 2024). Following this strategy, we consider average-, band depth-, energy score-, and dependence-based histograms in our analysis.
In addition, the reliability index (RI) provides a quantitative measure of the uniformity of a rank histogram. For a histogram with $K+1$ bins and predicted frequencies $\hat{p}_{k}$ in each bin, the reliability index is defined as
@@
Figure 2: Locations of solar irradiance observation stations in northern Chile (a) and visibility observation stations across Central Europe (b). Black lines indicate edges used to construct the input graph for a GraphSAGE model.
@@ 4 Implementation details
For visibility, each observation station is represented by a feature set that includes the control forecast, the mean and standard deviation of the 50 exchangeable ensemble members, and the proportions of forecasts predicting visibility below 5000 m, between 5000 and 30,000 m, and 30,000–70,000 m. Additionally, point forecasts from the Copernicus Atmosphere Monitoring Service (CAMS) are incorporated, along with annual base functions to capture seasonal and temporal effects, and the station’s latitude, longitude, and elevation. As with the solar irradiance data, this feature set largely follows Lakatos and Baran, (2024); however, for the GNN, the geographical coordinates are explicitly included to enable graph construction and capture location-dependent information. Lead time is also included to account for temporal variations.
@@
where $\alpha$ and $\beta$ are weighting coefficients. In our experiments, we set $w_{1}=1$ and $w_{2}=1-w_{1}$ for simplicity and interpretability, although in principle other values (e.g. $w_{1}>1$) could be used. Due to the large difference in scale between the two terms–the VS was found empirically to be several orders of magnitude larger than the ES–the VS component was normalized by an appropriate factor. Specifically, this factor was computed as the ratio of the mean ES to the mean VS over the raw ensemble, which was consistent with the ratios observed in the training batches. This normalization allowed us to select a weighting within the [01][0–1] interval that balances the contributions of both terms, ensuring that neither dominates the loss and that optimization proceeds stably. The GNN models were implemented in Python using the PyTorch Geometric library (Fey and Lenssen,, 2019).
@@ 5.2 Univariate performance of visibility forecasts
Figure 5 shows the mean CRPS of raw, POLR-, and GNN-based post-processed visibility forecasts, and the CRPSS of the post-processed forecasts relative to the POLR predictions, both as functions of lead time. As in the previous case study, Figure 5a indicates that all post-processed models outperform the raw ensemble forecasts; however, this advantage decreases with increasing lead time for all models. Differences in model performance are more clearly visible in Figure 5b, which presents the skill scores. Here, the GNN-CRPS model provides the largest improvement relative to POLR, followed by the dualGNN, while the GNN-ES model also shows positive gains over POLR for almost all lead times. The ranking of the GNN models is further confirmed by Table 3, which presents the overall mean CRPS of the post-processed visibility forecasts as a proportion of the mean CRPS of the raw ensemble, with the best-performing GNN-CRPS model showing a 2.02% improvement over POLR and the dualGNN showing a 1.23% improvement.
Finally, Figure 6 presents the coverage and average widths of the 90% central prediction intervals for raw and post-processed visibility forecasts as functions of lead time. Figure 6a shows the proportion of observations falling between the 5th and 95th percentiles of the forecasts. The GNN models exhibit a mean absolute deviation of approximately 1% from the nominal 90% coverage, compared to 2% for POLR and 51% for the raw ensemble. Moreover, temporal dependencies and the diurnal cycle are less pronounced in the coverage values of all post-processed models than in the raw ensemble.Figure 6b illustrates that the raw ensemble produces the narrowest prediction intervals, whereas the GNN models have narrower average widths than POLR and are less sensitive to the diurnal cycle.
@@ 5.3 Multivariate performance of solar irradiance forecasts
In this section, the performance of the proposed GNN models is assessed in comparison with a range of empirical copula-based approaches. Specifically, when the EMOS and MLP samples introduced in Section 5.1 are reordered to match the rank structure of the raw WRF forecasts, the resulting methods are denoted EMOS-ECC and MLP-ECC, respectively. Alternatively, when the dependence template is derived from past observations within the training period, the approaches are referred to as EMOS-SSh and MLP-SSh. As shown by Lakatos et al., (2023) restricting the set of possible pool dates to the training period does not deteriorate forecast performance. To provide a baseline, we also consider the EMOS and MLP samples under the assumption of independence.
Figure 7 displays the mean ES and the corresponding skill scores of the raw, independently post-processed, and multivariate post-processed solar irradiance forecasts as functions of lead time. As shown in Figure 7a, all post-processed models substantially improve the mean ES compared to the raw forecasts, and the ranking of the models remains consistent between the first and second 24 hours. Based on the mean ES, for cases when solar irradiance exceeds $7.5$ $W/m^{2}$, the models are ranked as follows: dualGNN, GNN-ES, GNN-CRPS, MLP-GNN, MLP-ECC, MLP-SSh, MLP, EMOS-SSh, EMOS-ECC, EMOS, and finally the raw ensemble forecasts. This ranking supports the conclusion that including the VS component in the loss function improves ES performance, indicating that incorporating VS into the training objective can enhance multivariate forecast evaluation based on ES. This is further shown in Figure 7b, where the ES-minimizing GNN is used as a reference. Based on this comparison, the EMOS samples exhibit the weakest performance relative to the reference, with EMOS-ECC and EMOS-SSh showing nearly identical skill. The same pattern holds for the subsequent MLP variants and the CRPS-trained GNN. Although MLP-ECC and MLP-SSh outperform the standard MLP forecasts and the GNN-CRPS, their performance is largely comparable. These models are followed by MLP-GNN reordered according to the dualGNN rank structure, which is surpassed only by dualGNN, yielding an average improvement of 4% over GNN-ES. The effect of including the VS in the loss function is most pronounced at the beginning and end of the day, i.e., during periods of low solar irradiance above 7.5 $W/m^{2}$, which correspond to the early morning and late afternoon hours. This suggests that the benefits of VS are particularly evident under low-irradiance conditions, where the spatial dependence captured by VS contributes most to improving forecast skill.
@@
Figure 9: Boxplots of Diebold–Mariano (DM) test statistics investigating the significance of the difference in ES (a) from the reference MLP-ECC and (b) GNN-ES methods as functions of the lead time. Grey region indicates the acceptance region of the two-tailed DM test for equal predictive performance at a 5% level of significance.
Figure 8, which shows the mean VS and corresponding skill scores of the raw, independently post-processed, and multivariate post-processed solar irradiance forecasts relative to GNN-ES, presents a very similar pattern. Based on Figure 8a, during periods with solar irradiance above $7.5$ $W/m^{2}$, the models are ranked as follows: dualGNN, GNN-ES, GNN-CRPS, MLP-GNN, MLP-SSh, MLP, MLP-ECC, EMOS-SSh, EMOS-ECC, EMOS, and finally the raw ensemble. Once again, the naive EMOS model exhibits the weakest performance, followed by EMOS-ECC and EMOS-SSh. The ranking of the EMOS variants is therefore identical to that observed in Figure 7a; however, the differences between models are more pronounced. Similarly to Figure 7a, the model ranking is consistent across the two consecutive days. Differences among the GNN and MLP models are more clearly visible in Figure 8b, where GNN-ES is again used as the reference. Relative to this benchmark, the EMOS model exhibits an average deficit of 65.82%, followed by EMOS-ECC with 43.32% and EMOS-SSh with 33.52%. In contrast, MLP and GNN-CRPS models show substantial improvements over the EMOS variants, with skill scores of -2.2%, -0.9%, and -0.8% for GNN-CRPS, MLP, and MLP-ECC, respectively. The MLP-SSh achieves a 0.1% improvement relative to GNN-ES, effectively matching the reference model’s performance. The model reordered according to the GNN rank structure shows a 1.3% advantage, while dualGNN achieves an average improvement of 5.2% over GNN-ES.
@@
Figure 10: Boxplots of Diebold–Mariano (DM) test statistics investigating the significance of the difference in VS (a) from the reference MLP-SSh and (b) GNN-ES methods as functions of the lead time. Grey region indicates the acceptance region of the two-tailed DM test for equal predictive performance at a 5% level of significance.
@@
As shown in Figure 10a, dualGNN again yields the largest gains over the strongest two-step benchmark, surpassing it in 76.92% of cases. The MLP-GNN ranks second, showing improvements in 65.38% of cases; notably, the benchmark model is never outperformed by either model. These forecasts are followed by GNN-ES (46.15%), MLP (7.69%), and MLP-ECC (3.85%). By contrast, GNN-CRPS and all EMOS variants never achieve significantly better performance than MLP-SSh; moreover, GNN-CRPS attains statistically indistinguishable skill in 92.31% of cases, whereas the EMOS models, in line with earlier findings, are consistently outperformed.
@@
Figure 11 presents the multivariate rank histograms of the post-processed forecasts, with the associated RIs indicated on the plots; the lowest RIs are highlighted in bold. While the RI provides a convenient summary of deviations from uniformity, interpreting multivariate rank histograms requires caution, as their meaning depends on the choice of pre-ranking method. In general, the two-step post-processing models and the GNN-based approaches produce the most uniform average- and band-depth rank histograms, with dualGNN exhibiting the most even distributions and the naive EMOS model the highest RI. All models show a left-skew in the Energy Score rank histograms, with the effect being least pronounced for GNN-ES. Based on the VS-based Dependence Rank histograms, the histogram corresponding to the MLP reordered according to the GNN rank structure is the most uniform, whereas deviations in the correlation structure would manifest as right- or left-skewed histograms.

### v2

@@ Abstract
The aim of this study is the multivariate post-processing of ensemble forecasts using a graph neural network (dualGNN) trained with a composite loss function that combines the energy score (ES) and the variogram score (VS). The method is evaluated on two datasets: WRF-based solar irradiance forecasts over northern Chile and ECMWF visibility forecasts for Central Europe. The dualGNN consistently outperforms all empirical copula-based post-processed forecasts and shows significant improvements compared to graph neural networks trained solely on either the continuous ranked probability score (CRPS) or the ES, according to the evaluated multivariate verification metrics. Furthermore, for the WRF forecasts, the rank-order structure of the dualGNN forecasts captures valuable dependency information, enabling a more effective restoration of spatial relationships than either the raw numerical weather prediction ensemble or historical observational rank structures. Notably, incorporating VS into the loss function improved the univariate performance for both target variables compared to training on ES alone. Moreover, for the visibility forecasts, the ES–VS combination even outperformed the strongest calibrated reference in terms of univariate performance.
@@ 3.2 Univariate post-processing of visibility
$P\!\big{(}Y\leq y_{i}\mid\mathbf{x}\big{)}=\frac{\exp\!\big{(}L_{i}(\mathbf{x})\big{)}}{1+\exp\!\big{(}L_{i}(\mathbf{x})\big{)}},\qquad L_{i}(\mathbf{x}):=\alpha_{i}+\mathbf{x}^{\top}\boldsymbol{\beta},\qquad i=1,2,\ldots,84,$
where $\alpha_{i}\in\mathbb{R}$ and $\boldsymbol{\beta}\in\mathbb{R}^{M}$ are model parameters satisfying $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{84}$. Consequently, fitting a POLR model to the visibility data requires estimating $84+M$ parameters. In the present work, we apply the local POLR model, as also used by Lakatos and Baran, (2024), for the calibration of visibility forecasts. In their study, this model provided the best predictive performance in terms of univariate performance across various spatial selection procedures, including semi-local and regional estimates (briefly discussed in Section 3.1), where local indicates that a separate model is fitted for each station based solely on its own historical data.
@@ 3.4 Verification metrics
For a predictive CDF $F$ and an observation $y\in\mathbb{R}$, the CRPS is defined as
@@
Moreover, for an ensemble forecast $\{f_{1},\dots,f_{K}\}$ resulting in an empirical CDF $\hat{F}_{K}$ the above reduces to the sample CRPS, that is
@@
The generalization of Equation (3.1) leads to the sample energy score (ES; Gneiting and Raftery,, 2007). For an ensemble forecast $\mathbf{f}_{k}=\big{(}f_{k}^{(1)},f_{k}^{(2)},\dots,f_{k}^{(D)}\big{)}^{\top}$, $k=1,\dots,K$, and the corresponding observation vector $\mathbf{y}=\big{(}y^{(1)},y^{(2)},\dots,y^{(D)}\big{)}^{\top}$, the sample energy score is given by
@@
Graphical tools for assessing the reliability of multivariate forecasts, such as multivariate rank histograms, are also well established. These histograms are based on pre-ranks, which condense a multivariate quantity into a single value. Since interpreting multivariate histograms can be challenging, it is advisable to employ multiple pre-ranking methods concurrently to obtain a comprehensive view of forecast miscalibration (Allen et al.,, 2024). Following this strategy, we consider average-, band depth-, energy score-, and dependence-based histograms in our analysis. The average rank is computed as the mean of the ranks of observations within the ensembles. Its histogram offers a diagnostic assessment of the forecast distribution: underdispersion is reflected by a $\cup$-shaped histogram, overdispersion by a $\cap$-shaped histogram, and bias typically results in a triangular pattern, while the band depth rank of each observation reflects its centrality relative to the ensemble of forecasts Thorarinsdottir et al., (2016). Whereas the energy score and dependence histograms build on the more recent concept of proper score-based pre-ranks introduced by Knüppel et al., (2022).
In addition, the reliability index (RI) provides a quantitative measure of the uniformity of a rank histogram, thereby indicating the degree to which the ensemble is properly calibrated. For a histogram with $K+1$ bins and predicted frequencies $\hat{p}_{k}$ in each bin, the reliability index is defined as
@@
Figure 2: Locations of solar irradiance observation stations in northern Chile (a) and visibility observation stations across Central Europe (b). Black lines indicate edges used to construct the input graph for the GraphSAGE models.
@@ 4 Implementation details
For visibility, each observation station is represented by a feature set that includes the control forecast, the mean and standard deviation of the 50 exchangeable ensemble members, and the proportions of forecasts predicting visibility below 5000 m, between 5000 and 30,000 m, and 30,000–70,000 m. Additionally, visibility point forecasts from the Copernicus Atmosphere Monitoring Service (CAMS) are incorporated, along with annual base functions to capture seasonal and temporal effects, defined as
$\beta_{1}(d):=\sin\big{(}2\pi d/365\big{)}\qquad\text{and}\qquad\beta_{2}(d):=\cos\big{(}2\pi d/365\big{)},$
where  $d$  denotes the day of the year. Furthermore, the station’s latitude, longitude, and elevation are also taken into account.
As with the solar irradiance data, this feature set largely follows Lakatos and Baran, (2024); however, for the GNN, the geographical coordinates are explicitly included to enable graph construction and capture location-dependent information. Lead time is also included to account for temporal variations.
@@
where $\alpha$ and $\beta$ are weighting coefficients. In our experiments, we set $w_{1}\in[0,1]$ and $w_{2}=1-w_{1}$, for simplicity and interpretability, although in principle other values (e.g. $w_{1}>1$) could be used. Due to the large difference in scale between the two terms–the VS was found empirically to be several orders of magnitude larger than the ES–the VS component was normalized by an appropriate factor. Specifically, this factor was computed as the ratio of the mean ES to the mean VS over the raw ensemble, which was consistent with the ratios observed in the training batches. This normalization allowed us to select a weighting within the $[0,1]$ interval that balances the contributions of both terms, ensuring that neither dominates the loss and that optimization proceeds stably. The GNN models were implemented in Python using the PyTorch Geometric library (Fey and Lenssen,, 2019).
@@ 5.2 Univariate performance of visibility forecasts
Figure 5 shows the mean CRPS of raw, POLR-, and GNN-based post-processed visibility forecasts, and the CRPSS of the post-processed forecasts relative to the POLR predictions, both as functions of lead time. As in the previous case study, Figure 5a indicates that all post-processed models outperform the raw ensemble forecasts; however, this advantage decreases with increasing lead time for all models. Differences in model performance are more clearly visible in Figure 5b, which presents the skill scores, with respect to the reference POLR model. Here, the GNN-CRPS model provides the largest improvement relative to POLR, followed by the dualGNN, while the GNN-ES model also shows positive gains over POLR for almost all lead times. The ranking of the GNN models is further confirmed by Table 3, which presents the overall mean CRPS of the post-processed visibility forecasts as a proportion of the mean CRPS of the raw ensemble, with the best-performing GNN-CRPS model showing a 2.02% improvement over POLR and the dualGNN showing a 1.23% improvement.
Finally, Figure 6 presents the coverage and average widths of the 90% central prediction intervals for raw and post-processed visibility forecasts as functions of lead time. The GNN models exhibit a mean absolute deviation of approximately 1% from the nominal 90% coverage, compared to 2% for POLR and 51% for the raw ensemble. Moreover, temporal dependencies and the diurnal cycle are less pronounced in the coverage values of all post-processed models than in the raw ensemble. Figure 6b illustrates that the raw ensemble produces the narrowest prediction intervals, whereas the GNN models have narrower average widths than POLR and are less sensitive to the diurnal cycle.
@@ 5.3 Multivariate performance of solar irradiance forecasts
In this section, the performance of the proposed GNN models is assessed in comparison with a range of empirical copula-based approaches. Specifically, when the EMOS and MLP samples introduced in Section 5.1 are reordered to match the rank structure of the raw WRF forecasts, the resulting methods are denoted EMOS-ECC and MLP-ECC, respectively. Alternatively, when the dependence template is derived from past observations within the training period, the approaches are referred to as EMOS-SSh and MLP-SSh. As shown by Lakatos et al., (2023) restricting the set of possible pool dates to the training period does not deteriorate forecast performance. Furthermore, for the current dataset, we include in the analysis a hybrid model that combines the MLP with the dualGNN by mapping the MLP-generated samples onto the rank order structure of the dualGNN forecasts for the same verification day and forecast horizon. Consequently, this hybrid model requires running the MLP and dualGNN in parallel. This reference model, hereafter referred to as MLP-GNN, enables us to evaluate whether the dualGNN rank structure provides additional information beyond that contained in the raw ensemble forecasts and historical observations. To provide a baseline, we also consider the EMOS and MLP samples under the assumption of independence, meaning that the calibrated samples are not rearranged according to any dependence template.
@@
Figure 9: Boxplots of Diebold–Mariano (DM) test statistics investigating the significance of the difference in ES (a) from the reference MLP-ECC and (b) GNN-ES methods as functions of the lead times 12–24 h and 36–48 h. Grey region indicates the acceptance region of the two-tailed DM test for equal predictive performance at a 5% level of significance.
Figure 7 displays the mean ES and the corresponding skill scores of the post-processed solar irradiance forecasts, both independent and multivariate, as functions of lead time, with the raw forecasts omitted from the skill score plot to better highlight the differences between the post-processed models. As shown in Figure 7a, all post-processed models substantially improve the mean ES compared to the raw forecasts, and the ranking of the models remains consistent between the first and second 24 hours. Based on the mean ES, for cases when solar irradiance exceeds $7.5$ $W/m^{2}$, the models are ranked as follows: dualGNN, GNN-ES, GNN-CRPS, MLP-GNN, MLP-ECC, MLP-SSh, MLP, EMOS-SSh, EMOS-ECC, EMOS, and finally the raw ensemble forecasts. This ranking supports the conclusion that including the VS component in the loss function improves ES performance, indicating that incorporating VS into the training objective can enhance multivariate forecast evaluation based on ES. This is further shown in Figure 7b, where the ES-minimizing GNN is used as a reference. Based on this comparison, the EMOS samples exhibit the weakest performance relative to the reference, with EMOS-ECC and EMOS-SSh showing nearly identical skill. The same pattern holds for the subsequent MLP variants and the CRPS-trained GNN. Although MLP-ECC and MLP-SSh outperform the standard MLP forecasts and the GNN-CRPS, their performance is largely comparable. These models are followed by MLP-GNN reordered according to the dualGNN rank structure, which is surpassed only by dualGNN, yielding an average improvement of 4% over GNN-ES. The effect of including the VS in the loss function is most pronounced at the beginning and end of the day, i.e., during periods of low solar irradiance above 7.5 $W/m^{2}$, which correspond to the early morning and late afternoon hours. This suggests that the benefits of VS are particularly evident under low-irradiance conditions, where the spatial dependence captured by VS contributes most to improving forecast skill.
@@
Figure 10: Boxplots of Diebold–Mariano (DM) test statistics investigating the significance of the difference in VS (a) from the reference MLP-SSh and (b) GNN-ES methods as functions of the lead times 12–24 h and 36–48 h. Grey region indicates the acceptance region of the two-tailed DM test for equal predictive performance at a 5% level of significance.
Figure 8, which shows the mean VS and corresponding skill scores of the raw, independently post-processed, and multivariate post-processed solar irradiance forecasts relative to GNN-ES, presents a very similar pattern. When forecasts are ranked according to the mean VS during periods in which the solar irradiance exceeds $7.5W/m^{2}$, the models are ordered as follows: dualGNN, GNN-ES, GNN-CRPS, MLP-GNN, MLP-SSh, MLP, MLP-ECC, EMOS-SSh, EMOS-ECC, EMOS, and finally the raw ensemble. Once again, the independent EMOS model exhibits the weakest performance, followed by EMOS-ECC and EMOS-SSh. The ranking of the EMOS variants is therefore identical to that observed in Figure 7a; however, the differences between models are more pronounced. Similarly to Figure 7a, the model ranking is consistent across the two consecutive days. Differences among the GNN and MLP models are more clearly visible in Figure 8b, where GNN-ES is again used as the reference. Relative to this benchmark, during periods with solar irradiance above $7.5$ $W/m^{2}$, the EMOS model exhibits an average deficit of 65.82%, followed by EMOS-ECC with 43.32% and EMOS-SSh with 33.52%. In contrast, MLP and GNN-CRPS models show substantial improvements over the EMOS variants, with skill scores of -2.2%, -0.9%, and -0.8% for GNN-CRPS, MLP, and MLP-ECC, respectively. The MLP-SSh achieves a 0.1% improvement relative to GNN-ES, effectively matching the reference model’s performance. The model reordered according to the GNN rank structure shows a 1.3% advantage, while dualGNN achieves an average improvement of 5.2% over GNN-ES.
@@
As shown in Figure 10a, which evaluates the significance of the VS deviations relative to the strongest GNN-independent MLP-SSh reference forecast, the dualGNN again yields the largest gains over the strongest two-step benchmark, surpassing it in 76.92% of cases. The MLP-GNN ranks second, showing improvements in 65.38% of cases; notably, the benchmark model is never outperformed by either model. These forecasts are followed by GNN-ES (46.15%), MLP (7.69%), and MLP-ECC (3.85%). By contrast, GNN-CRPS and all EMOS variants never achieve significantly better performance than MLP-SSh; moreover, GNN-CRPS attains statistically indistinguishable skill in 92.31% of cases, whereas the EMOS models, in line with earlier findings, are consistently outperformed.
@@
Figure 11 presents the multivariate rank histograms of the post-processed forecasts, with the associated RIs indicated on the plots; the lowest RIs are highlighted in bold. While the RI provides a convenient summary of deviations from uniformity, interpreting multivariate rank histograms requires caution, as their meaning depends on the choice of pre-ranking method. In general, the two-step post-processing models and the GNN-based approaches produce the most uniform average- and band-depth rank histograms, with dualGNN exhibiting the most even distributions and the independent EMOS model the highest RI. All models show a left-skew in the Energy Score rank histograms, with the effect being least pronounced for GNN-ES. Based on the VS-based Dependence Rank histograms, the histogram corresponding to the MLP reordered according to the GNN rank structure is the most uniform, whereas deviations in the correlation structure would manifest as right- or left-skewed histograms.